A 55.0 m wire carries a current of 2.25 A in a location where Earth’s


magnetic field is 4.5 x 10^-3 T. The wire experiences a force of 0.15 N. Find


the angle between the wire and the magnetic field.

1
Expert's answer
2022-04-06T15:44:16-0400

The magnetic field force on a wire with current

"F=BIl\sin\theta"

Hence

"\theta=\sin^{-1}\frac{F}{BIl}"

"\theta=\sin^{-1}\frac{0.15}{4.5*10^{-3}*2.25*55.0}=15.6^\circ"